# From Grounding to Manipulation: Case Studies of Foundation Model Integration in Embodied Robotic Systems

### v1

@@ 3 Case Studies on Instruction Grounding
$\boldsymbol{\pi_{0}}$finetuned 96.8 98.8 95.8 85.2 94.15
[…]
@@ 4 Case Studies on Robotic Manipulation
Given the scale of VLAs, we compare partial fine-tuning, which leverages our benchmark dataset (Appendix) and its inherent distribution bias to study convergence behavior, and full fine-tuning, which uses large-scale datasets to minimize the training loss. Our evaluation focuses on three key aspects: (i) training dynamics—how quickly and smoothly training converges; (ii) generalization—how well the resulting policies perform on various tasks; and (iii) robustness—how well the resulting policies handle environmental distractors. Our experiments highlight the performance of VLA models in different settings, offering practical suggestions for practitioners who have to adapt large VLAs under tight data, time and compute budgets.
Skill Adaptation Performance.
@@
Our fine-tuning process consists of two stages: (1) To assess convergence behavior under distribution shift, we collected a custom dataset (see Appendix for details) with a distribution bias relative to common pretraining datasets included in the Open-X-embodiment Collaboration et al. (2024) and LIBERO datasets Liu et al. (2023). We used it to partially fine-tune several recent VLA models and trained Diffusion Policy (DP) and Action Chunking Transformer (ACT) from scratch. The results are shown in Fig. 5; (2) For full fine-tuning, we leveraged larger benchmark datasets, Open-X-embodiment and LIBERO, to fully fine-tune RT-1, OpenVLA, SpatialVLA and NORA, and compared their performance. The results are shown in Fig. 6.
  • Partial Fine-tuning. Through the experiments we observe that DP and ACT exhibit high stability with low variance during training. In contrast, generalist models such as OpenVLA and $\pi_{0}$ require significantly more training iterations to attain comparable accuracy and exhibit greater variance, which can be attributed to their large model capacity. Notably, although DP achieves lower loss by fitting directly to noise, it still demands more training steps to generate coherent actions, even after loss convergence.
@@
  • Full Fine-tuning. These fully fine-tuned VLA models are evaluated on three tasks: (1) out-of-distribution (OOD) object manipulation, (2)  spatial relationship reasoning and (3) multi-object pick-and-place tasks. In task (1), both NORA and OpenVLA succeed, while SpatialVLA fails due to incorrect affordance point estimation. In task (2), NORA correctly follows instructions, while OpenVLA fails and SpatialVLA exhibits unstable performance. In task (3), NORA achieves successful execution while other models fail to complete the task reliably.
Sim2Real Adaption Performance.
We compare model performance on simulation benchmarks and real robot deployments. A significant drop in performance is observed during transfer from simulation to the real world (Table 3; Appendix Table 6). The simulation benchmark includes 30 procedurally-generated, disentangled tasks requiring nuanced spatial reasoning (LIBERO-Spatial), object understanding (LIBERO-Object), and goal interpretation (LIBERO-Goal), as well as 10 long-horizon, entangled tasks (LIBERO-Long).
@@
To evaluate robustness, we introduced distractor objects into the environment. As shown in Table 8, both OpenVLA and NORA exhibit substantial performance degradation in the presence of these perturbations, highlighting their sensitivity to novel conditions.
Figure 5: Results for partial fine-tuning of VLA models including OpenVLA and $\pi_{0}$, alongside results from training Diffusion Policy and ACT (Action Chunking Transformer) from scratch on our dataset. VLA models require more training epochs to converge and exhibit higher variance in performance.
@@
Current VLA models still face significant limitations in the following areas:
  • Adaptation and Generalization. A generic robotic policy is expected to quickly adapt to datasets with distributional shifts. However, according to the partial fine-tuning results, due to the large model capacities and the limited size of task-specific datasets, these VLA models failed to achieve fast adaptation. While full fine-tuning offers improved performance, it requires extensive data and long training time, which are impractical for many real-world scenarios.
  • Robustness. Robustness to distribution shifts (without finetuning) is a critical challenge. Results reveal substantial performance degradation both when encountering unseen objects and during sim-to-real transfer, highlighting the fragility of current VLA models in dynamic and unpredictable environments.
These findings suggest that while VLA models hold promise, they have limitations in data efficiency, adaptation speed, and robustness to make them reliable for real-world robotic applications.
Figure 6: Success rates of fully fine-tuned VLA models on out-of-distribution object manipulation (OOD Object), spatial relationship reasoning (Spatial) and multi-object pick-and-place (Multiple) tasks. NORA achieves the highest performance.
@@ 5 Constraints and Future Directions
Despite the promise of foundation models for enabling embodied agents to perform daily tasks, the following critical constraints still hinder their widespread deployment:
@@
In contrast to natural language datasets, which are readily sourced from internet, robotic datasets are significantly more expensive due to high hardware costs and intensive labor during data acquisition. A promising direction for future research is developing more data-efficient models. In addition, exploring high-fidelity simulation environments and developing robust sim-to-real transfer techniques could mitigate data scarcity.
Limited Generalization Capability.
A key limitation of current VLA models is their limited ability to generalize to out-of-distribution concepts that were not well-represented during training, which is a consequence of the aforementioned data scarcity. Many models depend heavily on large-scale paired datasets, which often exhibit biases and limited diversity in aspects such as camera viewpoints, lighting conditions and specific robotic embodiments. This results in fragile performance when deployed in real-world or domain-specific scenarios. Furthermore, these models struggle with fine-grained spatial reasoning and temporal understanding, hindering them from accurately aligning language with complex visual scenes or dynamic events.
@@
Deploying large-scale models on robotic platforms introduces significant computational challenges, primarily in inference speed and GPU RAMs. This underscores the importance of smaller models that can efficiently generate actions without significant performance degradation. This issue is particularly pronounced in autoregressive models, and diffusion models are less affected.
@@
When evaluating the generalization capabilities of Vision-Language Alignment (VLA) models, this paper primarily focuses on different tasks without extensively addressing their generalization across varying robot morphologies. Although relatively few works have specifically targeted this aspect, it remains a significant challenge in deploying VLA models in real-world applications. Robots with different morphologies, such as bimanual manipulators, humanoid robots, and autonomous vehicles, require distinct operational protocols and safety considerations. The absence of a generic robot policy that can adapt seamlessly across diverse morphologies limits the practical generalization potential of VLA models and hinders their deployment as universal robotic policies.
In addition, this paper does not explore the ability of VLA models to ground instructions involving open-ended or ambiguous commands. Current VLA models are largely trained on curated datasets, which allow them to learn mappings from instructions to specific actions. However, this reliance constrains their ability to truly understand instructions at the semantic level. As a result, when facing out-of-distribution or vague commands, these models often struggle to infer reasonable actions. Addressing this limitation will require integrating more advanced instruction-understanding modules into the VLA pipeline to improve their robustness in handling ambiguous or under-specified input.
@@ A.2 Complex Instruction Grounding Dataset
We curated an evaluation dataset for the complex instruction grounding task in cluttered scenes [^4] https://github.com/xiuchao/InstructionGrounding. Thirty images were sampled from the action sequences and subsequently categorized based on the number of objects: Easy (<15), Medium (15$\sim$20), and Hard (>20). Objects in the visual scenes were manually annotated using visual prompts and paired with various instructions. The spatial relationship words were illustrate in Table 4.
@@ C.1 Fine-tuning details for VLA
Full fine-tuning of RT-1, OpenVLA, SpatialVLA, and NORA was conducted on a compute node equipped with 8$\times$×H100 GPUs. The fine-tuned models were evaluated on 9 diverse real-world manipulation tasks, as shown in Fig. 11. Success rates are summarized in Table 6, demonstrating NORA’s superior policy generation capabilities across three task categories: out-of-distribution object grasping, spatial reasoning, and multi-object manipulation.
@@ C.2 Impact of Action Chunking
Subsequently, we modified the execution policy to only perform the first action in each predicted chunk. This adjustment resolved the collision issue and NORA -LONG achieved an 80% success rate on the “put the carrot in the pot” task. However, on multi-object pick-and-place tasks, NORA-LONG consistently stopped after placing the first object, resulting in a 0% final success rate. For the spatial reasoning task, NORA-LONG achieved a 70% success rate on “put the pink toy at the right corner.”

### v2

@@ 3 Case Studies on Instruction Grounding
$\boldsymbol{\pi_{0}}\$finetuned 96.8 98.8 95.8 85.2 94.15
$\pi_{0}\text{-FAST}$ finetuned 96.4 96.8 88.6 60.2 85.5
@@ 4 Case Studies on Robotic Manipulation
Given the scale of VLAs, we compare partial fine-tuning, which leverages our curated probing dataset (Fig. 3) and its inherent distribution bias to study convergence behavior, and full fine-tuning, which uses large-scale datasets to minimize the training loss. Our evaluation focuses on three key aspects: (i) training dynamics—how quickly and smoothly training converges; (ii) generalization—how well the resulting policies perform on various tasks; and (iii) robustness—how well the resulting policies handle environmental distractors. Our experiments highlight the performance of VLA models in different settings, offering practical suggestions for practitioners who have to adapt large VLAs under tight data, time and compute budgets.
Real-World Skill Adaptation.
@@
Our fine-tuning process consists of two stages: (1) To assess convergence on tasks beyond the coverage of common pretraining datasets, we collected a custom dataset (Appendix A.1) focused on screwdriver-picking in cluttered tabletop scenes. This task introduces both object- and scene-level distribution gaps, as screwdrivers and dense clutter are largely absent from Open-X-Embodiment Collaboration et al. (2024). We used this dataset to partially fine-tune generalist VLA models, and to train compact task-specific models such as Diffusion Policy (DP) and Action Chunking Transformer (ACT) from scratch. (2) For full fine-tuning, we leveraged larger benchmarks—Open-X-Embodiment and the simulated LIBERO dataset Liu et al. (2023)—to fully fine-tune RT-1, OpenVLA, SpatialVLA, and NORA, and compared their performance.
Figure 5: Partial fine-tuning results for VLAs (OpenVLA and $\pi_{0}$) compared with training Diffusion Policy (DP) and ACT from scratch on our dataset. VLAs require more epochs to converge and show higher performance variance.
Figure 6: Success rates of fully fine-tuned VLAs on multi-object pick-and-place, out-of-distribution object manipulation, and spatial relationship reasoning tasks. NORA achieves the highest performance.
  • Partial Fine-tuning. We observe that DP and ACT converge stably with low training variance (Fig. 5). In contrast, generalist models such as OpenVLA and $\pi_{0}$ require far more iterations to reach comparable accuracy and exhibit greater variance. This difference reflects model scale: compact task-specific policies like DP and ACT contain only tens of millions of parameters, whereas generalist policies such as OpenVLA (7B) and $\pi_{0}$ (3B) rely on billion-scale backones, making optimization slower and less stable. Notably, while DP attains lower loss by fitting directly to noise, it still requires additional training to produce coherent action sequences even after loss convergence.
@@
  • Full Fine-tuning. The fine-tuned VLAs are evaluated on three tasks: (1) out-of-distribution object manipulation, (2) spatial relationship reasoning, and (3) multi-object pick-and-place. In task (1), both NORA and OpenVLA succeed, whereas SpatialVLA fails due to incorrect affordance-point estimation. In task (2), NORA follows instructions correctly, while OpenVLA fails and SpatialVLA shows unstable performance. In task (3), only NORA executes the task successfully, while the other models fail to complete it reliably (Fig. 6).
Simulated Skill Adaptation.
We compare model performance on simulation benchmarks and real robot deployments (Table 3). The results show that the finetuned $\pi_{0}$ model outperforms all baselines across tasks, achieving the highest success rates on Spatial and Object tasks while maintaining strong performance on Goal and Long-horizon tasks. Furthermore, the ablation study of action chunking (AC) on NORA demonstrates that AC consistently enhances performance across most simulation tasks. Notably, a significant drop in performance is observed when transferring from simulation to the real world (Appendix Table 6).
@@
We assess robustness by introducing distractor objects into the environment. As shown in Appendix Table 8, both OpenVLA and NORA degrade substantially under these perturbations, underscoring their sensitivity to novel conditions.
@@
Current VLAs still face significant limitations in the following areas:
  • Adaptation and Generalization. A generic robotic policy is expected to adapt rapidly to datasets with distributional shifts. However, our partial fine-tuning results show that, given their large model capacities and the limited size of task-specific datasets, current VLAs fail to achieve efficient adaptation. While fine-tuning improves performance, it demands extensive data and prolonged training, which is usually impractical for many real-world scenarios.
  • Robustness. Robustness to distribution shifts without finetuning remains a critical challenge. Results reveal substantial degradation when encountering unseen objects and during sim-to-real transfer, underscoring the fragility of current VLA models in dynamic and unpredictable environments.
These findings indicate that although VLAs hold significant promise, they remain constrained by poor data efficiency, slow adaptation, and brittle robustness. Bridging these gaps will require both algorithmic advances—such as more parameter-efficient adaptation, bias-resistant pretraining, and stronger perception backbones—and system-level improvements in data collection and augmentation. Addressing these limitations is crucial if VLAs are to evolve from research prototypes into reliable, deployable policies for real-world robots operating under uncertainty.
@@ 5 Constraints and Future Directions
Despite the promise of foundation models (FMs) for enabling embodied agents to perform daily tasks, several critical constraints still hinder their reliable deployment:
@@
Unlike natural language data that abundantly available on the internet, robotic datasets are costly to collect due to hardware wear, safety risks, and labor-intensive demonstrations. A key direction is improving data efficiency through parameter-efficient adaptation, imitation from unlabeled interaction, and self-supervised pretraining. Complementary approaches include leveraging high-fidelity simulators and developing robust sim-to-real transfer pipelines to reduce reliance on large real-world collections.
@@
VLAs place heavy computational burdens on robotic platforms, leading to bottlenecks in inference speed. The constraints motivate research on lightweight architectures and efficient decoding strategies that can sustain performance while satisfying the real-time requirements of embodied control.
Explainability and Safety.
Most FMs lack explicit mechanisms for interpretability or safety guarantees—factors that are crucial for deployment in high-stakes or unstructured environments. These models may output confident but incorrect actions when faced with out-of-distribution inputs or adversarial perturbations. Moreover, without built-in constraints to enforce ethical and operational boundaries, VLAs risk misinterpreting ambiguous instructions in ways that compromise human intent or physical safety.
@@
For VLA generalization study, this work focuses primarily on task-level performance and does not extensively examine generalization across diverse robot morphologies. Yet this remains a critical challenge for real-world deployment: robots with different embodiments—such as bimanual manipulators, humanoid robots, or mobile platforms—require distinct control protocols and safety constraints. The absence of a generic policy that adapts seamlessly across such morphologies limits the universality of current VLA approaches.
This work also does not directly investigate the grounding of instructions that are open-ended or ambiguous. Existing VLAs are trained largely on curated datasets that map well-structured commands to specific actions, but this reliance constrains their semantic understanding. Consequently, when faced with vague or out-of-distribution instructions, they often fail to infer reasonable behaviors. Addressing this limitation will require integrating richer language-understanding modules and more diverse training data to improve robustness in handling underspecified or ambiguous input.
Acknowledgment
This work is supported by the National Robotics Programme (Award #M23NKBK0091, #M25N4N2009). This work is also supported by the National Research Foundation, Singapore under its National Large Language Models Funding Initiative (AISG Award No: AISG-NMLP-2024-005), NTU SUG project #025628-00001:Post-training to Improve Embodied AI Agents, and A*STAR SERC CRF funding to Cheston Tan. We thank Wei Qi Toh and Ray Prithviraj for their assistance with robot experiments, and Shaohua Li for manuscript polishing.
@@ A.2 Complex Instruction Grounding Dataset
[…]
@@ C.1 Fine-tuning details for VLA
[…]
@@ C.2 Impact of Action Chunking
Subsequently, we modified the execution policy to only perform the first action in each predicted chunk. This adjustment resolved the collision issue but the model performance is still degraded.